The following selected transactions occurred during 2016 and 2017 for Baltic Importers. The company ends its accounting yearthe numbers in the 3rd box is wrong, unsure how to compute
Answer #1

Interest accrues on 1st note for 3 months as on September end

Interest on 1st note = 13000 * 0.08 * 3/12 = 260

Interest accrues on 2nd note for 24 days as on September end

Interest on 2nd note = 6000 * 0.12 * 24/360 = 48

Total interest accrued = 260 + 48 = 308

Journal:

Interest Receivable 308

Interest Revenue 308

(Being Interest accrued)